SOHO DAILY REPORT Operations Day - 043/2004 (12 February 04) Mission Day 2995 SOHO FLIGHT OPERATIONS TEAM, NASA/GSFC Ref. #: SH-FOT-DR-040212 ==================================================================== PASS #1 Canberra DSS46 0133-0855 PASS #2 Madrid DSS66 0845-1847 PASS #3 Goldstone DSS27 1839-2059 PASS #4 Goldstone DSS16 2039-044/0235 ********************************************************************* PASS #1 Canberra DSS46 0133-0855 Anomalies: G04-0062 0141 Errant lock on spacecraft receiver 2 due to incorrect uplink frequency setting - delayed AOS by 4 minutes. DR# C103328 G04-0063 0145 Ranging not performed - cause unknown - unable to collect 435 minutes of ranging data, delayed commanding & pass activities 54 minutes, lost 3 minutes of VC3 which was recovered from DSS27. DR# C103328 G04-0064 0145 Degraded telemetry - cause unknown - degraded data for 275 minutes - recoverable from CDR. DR# N101926 NRT: None Activities Summary: - Collected MDI-M (0717-0833) - Collected MDI-H (0133-0717, 0833-0834) - Collected Ranging Data - OBT Distribution to SWAN, CELIAS - SSR SEF Reset (2x) - SSR Record ------------------ PASS #2 Madrid DSS66 0845-1847 Anomalies: G04-0066 1358 Spacecraft receiver 2 lost lock - cause unknown - unable to command for 8 minutes, delayed MDI transition by 8 minutes. DR# M102296 G04-0065 1403 Degraded telemetry/lost telemetry due to possible RFI - lost 1.25 minutes of degraded telemetry, delayed an activity by 5 minutes. DR# M102295 NRT: 1449-1834 Activities Summary: - Collected MDI-M (0923-0942, 1408-1835) - Collected MDI-H (0942-1408, 1835-1847) - Reaction Wheel Speed Monitoring Daily Adjustment - CRS Drift Estimation Update - SSR SEF Reset (1x) - HGA Table Upload - Hx Limit Update - CAE-B in the AOCS HW Config Word ------------------ PASS #3 Goldstone DSS27 1839-2059 Anomalies: None NRT: 1907-2030 Activities Summary: - Collected MDI-M (1906-2035) - Collected MDI-H (1839-1906, 2035-2059) ------------------ PASS #4 Goldstone DSS16 2039-044/0235 Anomalies: None NRT: 2116-2243 Activities Summary: - Collected MDI-M (2115-044/0200) - Collected MDI-H (2039-2115, 044/0200-044/0235) - SSR SEF Reset (1x) ------------------ -- End of Report --